25 February 2008

RHIC Running:  Power dip on Sunday brought down the fill at 10:36AM, taking most of the day to recover.  Beam was not brought back in RHIC until after midnight this morning.  Blue abort kicker timing was wrong, late by 200ns, which made for dirty dumps.  To compensate for this, the first 4 bunches are filled only to low intensity.   

On Saturday night we went back to measuring the JET polarization in the Blue ring at about 9:00PM.  On Sunday the JET holding field magnet tripped off (again).  No JET measurements can be made until they get a short access to reset the cooling loop fault.

Polarization is low in both rings, <40% at store. Blue Vertical Target #4 broke, so we had to switch to the thick Blue Vertical target #5 for polarization measurements.  On Friday we lost Yellow Vertical Target #1.  We are losing one target a week this year.  Last year we were losing one target every two weeks.  Statistics are poor, but things do not look good for the remaining two weeks of the run.

Today there will be work on fixing the timing of the Blue abort kicker.  And continuing attempts to improve the polarization and intensity.

 


Scheduling Physicist: Mike Sivertz Last updated:
Email: sivertz@bnl.gov
Office Phone: 631-344-6102
Digital Pager: 7528 or 4300
Home Phone: 631-878-9020